OpenStudy (anonymous):

459 is 85% of what number?

OpenStudy (anonymous):

You should be able to try this by yourself... but once again... is = = %= x/100 what number= x of= times see if you can plug in these values.

OpenStudy (anonymous):

\[459 = {85\over100} \times x\]